### v2

# Communicating the gravitational-wave discoveries of the LIGO–Virgo–KAGRA Collaboration

Abstract
The LIGO–Virgo–KAGRA (LVK) Collaboration has made breakthrough discoveries in gravitational-wave astronomy, a new field of astronomy that provides a different means of observing our Universe. Gravitational-wave discoveries are possible thanks to the work of thousands of people from across the globe working together. In this article, we discuss the range of engagement activities used to communicate LVK gravitational-wave discoveries and the stories of the people behind the science using the activities surrounding the release of the third Gravitational Wave Transient Catalog as a case study.
Keywords: Popularization of science and technology; Public engagement with science and technology.
1 Introduction: Context and Objectives
Gravitational waves (GWs) can be difficult to imagine. They are ripples in spacetime, created by the acceleration of massive objects, that propagate across the Universe at the speed of light. The strongest GW sources are massive and rapidly moving: the best sources for ground-based GW detectors are coalescing binaries of black holes or neutron stars. After travelling across the cosmos to Earth, GWs are almost imperceptibly small. The strongest signals correspond to a stretching and squeezing of space of $1$ part in $10^{21}$. Observing GWs has therefore been a great experimental challenge, requiring an international community to design and build the complex, highly sensitive instruments needed to measure the waves, and to create and test the algorithms needed to analyse the data.
In 2015, one hundred years after Einstein first calculated the properties of GWs in his theory of general relativity, the first direct observation of GWs was made [1]. This discovery was confirmation for Einstein’s theory, an experimental triumph, and the beginning of a new era for astronomy. The signal GW150914 came from two black holes, each about $30$ times the mass of our Sun, coalescing more than a billion light-years away. This was the first time such a binary black hole system was found, the first time black holes of this size were discovered, and the first time two black holes were observed to inspiral and merge. The discovery was a major news event around the world [2], building on the public’s enduring fascination with Einstein and black holes.
Since 2015, the LIGO–Virgo–KAGRA (LVK) Collaboration, which operates the international network of GW observatories (the two LIGO sites in the US, Virgo in Italy, KAGRA in Japan and GEO 600 in Germany) has made many further GW discoveries. The field has grown rapidly, with the third GW Transient Catalog (GWTC-3) increasing the number of probable detections to $90$ [3]. The latest observing run is ongoing. These observations have revealed a diverse range of black hole and neutron star binaries, and revolutionised our understanding of the astrophysics of these sources.
Surrounding each results release are numerous engagement and communication activities. These share the wonder of discovery, what we can learn with GWs, and the technological advances that enable these groundbreaking measurements. Engagement and communication activities are carried out through the effort and enthusiasm of individual scientists, collaboration-wide teams, and partnerships between scientists and non-scientists, all of which complements the work of professional science communicators at the observatories. Matching the global composition of the LVK, selected resources have been translated into over $20$ languages.
Discoveries also provide an opportunity to focus on the people behind the science. The LVK includes around $2000$ people from diverse backgrounds across the globe; from designing instrumentation to operating the kilometre-scale laser interferometers, from coding analysis algorithms to calculating GW emission. By making connections with the people within the LVK, we show the human side of how science works, and provide possible role models for young people interested in science and technology.
In this article, we detail communication methods used by the LVK. We review the range of engagement activities used to communicate LVK GW discoveries and the stories of the people behind the science, with specific examples for GWTC-3.
2 General communication methods
Discovery announcements from the LVK normally have three components: paper, data release, and educational and engagement resources. Here we describe some general and long-term education and engagement activities of the LVK.
2.1 The people behind the science
Figure 1: Frequency of Humans of LIGO participants’ institution (top) and hometown (bottom) county. Scientists come from more diverse locations than just the institutions of the LVK.
Discovery science is not just about the results, it is also about the people behind them. Highlighting individuals within the LVK can provide STEM role models, challenge stereotypes, and help to foster a welcoming environment.
Figure 2: Antimatter comic showing LIGO scientists performing detector checks (credit: Nutsinee Kijbunchoo).
The LIGO Magazine (www.ligo.org/magazine) is published twice annually. Content is commissioned and edited by volunteers from the LIGO Scientific Collaboration and it has a professional, high-quality design for print and online PDF. Since the first edition in 2012, the LIGO Magazine has featured $470$ authors, representing $237$ workplaces in $23$ countries. Authors are from the GW community and beyond, and the target audience is both collaboration members and the interested public. Articles include perspectives of working on GW science, personal stories, opinion pieces, and advice columns.
Humans of LIGO (humansofligo.blogspot.com) showcases the lives, backgrounds, and inspirations of LVK members and GW scientists through blog posts. Each post consists of an image of the participant, a direct quote about their experiences, and a short biography including job description and hobbies. Since its inception in 2018, the blog has featured $59$ participants from $11$ different counties, working at $42$ institutions (Figure 1). The blog posts are shared on LIGO social media, and have reached over $53,000$ views (as of May 2024).
Antimatter web comics (antimatterwebcomics.com) by Nutsinee Kijbunchoo features day-to-day experiences and current affairs from the GW community and beyond (Figure 2). Mental health and graduate-student life are focuses of many comics. Comics have been regularly featured in the LIGO Magazine.
The LIGO-India blog, Glorious Women [4] showcases STEM role models, highlighting women from across the LVK. Behind-the-scenes interviews feature the daily lives of students, their motivation, contributions and challenges [5].
2.2 Academic audiences
Sharing our discoveries, and how to use them, is key to ensuring a wider impact on the scientific community.
Communication with academics is primarily through scholarly articles. Since GW astronomy is a new field, the LVK have written dedicated articles on introductory GW data analysis [6], the structure of our data releases [7, 8], and the basic physics of a binary black hole coalescence (suitable for undergraduate teaching) [9].
Data releases are coordinated by the GW Open Science Center (GWOSC), which hosts raw data, links to LVK publications and further data products (gwosc.org). GWOSC maintains lists of open-source analysis software and provides teaching on data analysis through annual Open Data Workshops. The workshops are delivered in a hybrid format; with 200200200200300300300300 people participating in-person each year, and over $7700$ people enrolling online over the last four years. Workshop materials are openly available online.
To accompany the release of key papers, the LVK organises webinars through Zoom, with attendees able to submit questions. Recordings are uploaded to YouTube.
The LVK also engages with the academic community to promote proper recognition of its three constituent collaborations. There is currently a bias in scholarly literature: an abbreviated narrative attributes credit to LIGO, excluding Virgo and KAGRA from LVK achievements, with consequences for scientific careers, accuracy of public media and interactions with funding agencies. A year-long project [10] studied this cognitive bias, and tried to educate the community.
2.3 Formal classroom education
LVK members have maintained long-term efforts to introduce GWs and related topics into formal classroom education. Examples include an Educator’s Guide [11], courses for community college teachers, high school resources, and undergraduate lab demonstrations [12].
Einstein-First (www.einsteinian physics.com) has developed an eight-year science curriculum that introduces quantum concepts and general relativity at primary school, followed by GWs, black holes, cosmology and quantum measurement in middle school [13]. Efforts are underway to introduce Einstein-First in Australia, India, USA and Brazil.
2.4 Science Summaries and press releases
LVK Science Summaries are a long-standing effort to explain LVK publications at a technical level typically higher than in the popular press but still accessible for enthusiastic lay readers. They are published on www.ligo.org, distributed as print copies at outreach events, and shared with journalists. Summaries are translated into a variety of languages by LVK volunteers, with discovery papers often receiving the most attention (e.g., the GW190521 summary was translated into 16 languages) [14].
Press releases to accompany discovery announcements are prepared centrally, including quotations from a selection of LVK members, and then adapted by member institutions (often including quotations from their scientists). These are frequently drawn upon for news articles.
2.5 In-person interactions: Science festivals, museums and visitor centers
In-person outreach to the general public reaches fewer people than other means, but it can yield impactful interactions, and may be perceived as having more value [15].
LVK members participate in science festivals, classroom visits, the creation of science exhibits (such as interactive detector models [16, 12], visitor centers and tours of the GW observatories. The face-to-face format allows staff to customize the experience to the visitor.
LVK’s visitor centers merge artifacts from the detectors with interactive exhibits that put the visitor in the role of a scientist or engineer. The exhibits explore the underlying science and engineering of GW detectors, from the importance of pendula to the concepts underpinning gravity. In-person and virtual tours of the detectors give a glimpse to the heart of operations. Over $330,000$ visitors have interacted with the LVK visitor centers; 2024 attendance is projected to reach $30,000$.
Science festivals and local outreach efforts around the world connect with classrooms and the general public. Outreach is performed by seasoned scientists, engineers, graduate students and undergraduates who have only recently learned about GWs. Such outreach efforts positively impact the undergraduates who perform the outreach [17, 18, 19].
2.6 Graphics
Figure 3: The detector data for GW150914, and an artistic representation of its binary black hole source (credit: LIGO/National Science Foundation/Aurore Simmonet/Somona State University). The source is shown at three stages, corresponding to the portions of the signal beneath.
Unlike electromagnetic observatories, GW data does not naturally lend itself to the stunning images that are commonplace throughout astronomy. Nevertheless, visualisations, infographics and cartoons [20]form key parts of our outreach activities, e.g., infographics displaying key facts are produced to accompany discoveries; simplified plots of results may be created for Science Summaries, and artistic representations of our sources may accompany press releases.
Videos and images are produced from numerical-relativity simulations of sources and the GWs they emit. Since new simulations are often produced to better understand the physics of novel sources, fresh visualisations have often accompanied discovery announcements.
Images accompanying our discovery announcements have been selected as NASA Astronomy Pictures of the Day, e.g., to accompany our first detection, an artistic representation of a binary black hole coalescence (Figure 3) and a corresponding numerical-relativity simulation were featured (11 and 12 February 2016), and to accompany GW190521, an artistic representation of the GW emission from two spinning black holes was featured (8 September 2020).
2.7 Multisensory resources: sound and touch
Sound is a common analogy when describing GWs. While GWs are not sound, the GW frequencies observable by the LVK are similar to audio frequencies of human hearing. The signals lend themselves to audification (www.soundsofspacetime.org): the technical name for the signal from a binary inspiral is a “chirp” in reference to its sweep up in frequency.
GWs can also be communicated through touch. The tidal spatial deformations of the first detected GW signal have been translated to a 3D-printable model, specifically tailored to the needs of visually impaired people. By running fingers along the edges of the shape, one can appreciate the features of a GW and perceive the details of the signal evolving in time. This design is freely available [21], along with suggestions for classroom activities.
Figure 4: A Tactile Universe 3D printed surface of the inspiral of two black holes and the resulting GWs (credit: Coleman Krawczyk).
The Tactile Universe (www.tactileuniverse.org[22] has partnered with UK LVK groups to design tactile resource sets for use with blind and vision-impaired pupils in upper secondary education. A mix of 3D-printable (Figure 4) and more basic tactile resources, paired with existing sonifications, are tied to 20 minute workshops that teach students about black holes and neutron stars, GW signals and detection, and core science topics like waves and gravity. The 3D-printable resources, example lesson scripts and guides are available on the website.
2.8 Art and music with science
Numerous art, music, and science collaborations explore novel approaches and interpretations of GWs. These collaborations showcase our science through a different and interesting point of view and often bring the topic of GWs to new audiences. Works have been created by both professional artists and musicians and members of the LVK [23].
Just a few examples include music compositions [24, 25], open land art installations [26], museum exhibits [27] and novel musical instruments (gravitysynth.tumblr.com). The LIGO-India blog has a virtual gallery GWSciArt that showcases GW-inspired artistic projects [28].
2.9 Social media and non-traditional media
Non-traditional media provides an opportunity to present our science to broad audiences. Activities like social-media posts, blogs, and podcasts or interviews can help communicate science in an informal way, e.g., the Spanish podcasts “Oscilador armónico”  [29], “A ciencia cierta” [30] and “Coffee Break: Señal y Ruido” [31] which coined and popularized the term “gravitondas” (“gravit-waves”) for GWs.
The LVK has several social media accounts, for the collaborations and the observatory sites, on X, Facebook, Instagram, Mastodon (the Astrodon instance). Content includes educational resources, news stories, and discovery announcements. Special posts are scheduled for events such as detection anniversaries or the International Day of Women and Girls in Science. Reddit has been used for Ask-me-anything posts about discoveries. Social media allows the LVK to interact directly with the public, and threads in response to questions have proved popular.
2.10 Computer games/apps, visualisations, and interactive technology
Interactive apps and games are engaging and fun ways to convey GW science. Physical interaction with virtual environments incorporates additional modes for audiences to engage with content. Activities range from computer games [32] (www.laserlabs.org/; blackholehunter.org) to classroom and science festival teaching with virtual reality [33, 34] (www.scivr.com.au).
3 Case Study: GWTC-3
GW discoveries from the LVK are published in catalogs: papers accompanied by data releases. GWTC-3 [3] is the most recent of these, presenting results up to the end of our third observing run. Thanks to the continued improvements in detector sensitivity [35], the rate of discoveries has increased with time, meaning that each catalog includes a significant number of new GW candidates. GWTC-3 represents the most sophisticated and comprehensive analysis published by the LVK to date.
Production and publication of a catalog is a multi-year project directly involving hundreds of experts. Given the significance of the results and the importance of easy-to-use data releases, effective communication is a priority. However, communication of catalog results faces: catalog results consist of comprehensive analyses of many signals (making them overwhelming), and the majority of detections have properties similar to those observed in the past (lacking the excitement of a novel discovery). For GWTC-3, the project team included a member responsible for coordinating communication and outreach activities from the beginning. This embedding allowed for these activities to be planned as the project progressed, developed in tandem with the data releases, and coordinated with those compiling the results.
The variety of activities designed to communicate the discoveries from GWTC-3 span the diverse range of audiences targeted and platforms used by the LVK.
3.1 Academic
A series of LVK webinars were organised to coincide with the release of GWTC-3. The first on GWTC-2.1 [36], a reanalysis of the previous data using updated methods consistent with GWTC-3. Subsequent webinars in the series presented the GWTC-3 results and implications for cosmology, astrophysics and tests of general relativity. Presenters were drawn from those who made key contributions to the work, with preference given to early-career scientists, and ensuring a balance in geographic location and demographics.
The data release was a key component of the GWTC-3 results. It included Jupyter notebooks and Python scripts to demonstrate use and how to reproduce plots from the paper. Additionally, a Streamlit app was created to interactively make plots of source parameters: this enabled users to customise paper plots without needing to download data (gwtc3-contours.streamlit.app/).
3.2 Written resources
Like the paper itself, the GWTC-3 Science Summary [37] had a broad scope to cover. It was written in close collaboration between outreach experts and the analysis and paper-writing teams. It includes background information needed to understand the catalog, and a subset of highlight discoveries. The summary has been translated into $9$ languages.
3.3 Magazine article
LIGO Magazine issue $20$ featured an $8$ page article on the story behind the catalog [38]. Short snippets from $23$ members of the analyses, paper writing, and engagement teams wrote about their experiences and stories of working on the catalog and associated results. Although many individuals contributed to this feature, they represent only a fraction of those providing input to GWTC-3. Articles like this give readers a behind-the-scenes insight of working on big LVK results.
3.4 Graphics and visualisations
GWTC-3 updates to some of our long-standing visualisations were made. The new results were added to the Masses in the Stellar Graveyard plot  5), showing the masses of black holes and neutron stars observed by the LVK alongside those observed via traditional electromagnetic astronomy. An online version allows for customisable images [39]. The LVK Orrery was also updated, which shows a stylised animation of the GWTC-3 binaries.
New visualisations were also created. One was a poster reminiscent of a periodic table showing the catalog after each observing period. The poster was used in a variety of articles, a double-page LIGO Magazine spread [40], and in the GWTC-3 webinar. An artistic representation of the detections in a box accompanied the publication in Physical Review X. An image showing spectrograms for the 90 GWTC-3 signals was featured as NASA Astronomy Picture of the Day (7 December 2021).
Figure 5: GWTC-3 version of Masses in the Stellar Graveyard (credit: LVK/A Geller/Northwestern). The plot shows the known masses of stellar-mass black holes and neutron stars observed electromagnetically and with GWs.
3.5 Audification
Low Mass Beats is a audification of the catalog, converting masses into pitch for each black hole or neutron star: the greater the mass, the lower the pitch. This lighthearted project provides an entertaining perspective on the catalog (SoundCloud: soundcloud.com/user-889003031).
3.6 Social media
GWTC-3 material was shared via a social media campaign in two bursts using the hashtag #GWCT3. The first burst coincided with the release of the preprint and associated data. Special attention was given to advertising the webinars and data releases. The second burst coincided with the publication of the journal article and priority was given to any resources developed since the initial release, such as the Streamlit plotter and Low Mass Beats.
For both bursts, posts were spaced over multiple days to catch the attention of people in different timezones. Material was translated into regional languages [41]. The pacing also made it natural to integrate popular-science articles written immediately after the release.
Resources were subsequently reshared, e.g., Low Mass Beats was shared on Halloween because of its spooky sound, and the Orrery was shared during NASA’s Black Hole Week.
The impact of the social media campaign can be partially assessed via the Altmetric score. As of June 2024, the GWTC-3 paper has a score of $713$ (including $56$ news outlets and 5 Wikipedia pages). This score is in the $99$th percentile of all tracked outputs.
4 Conclusions
GW astronomy is a new branch of astronomy. After decades of research, the first observation was made in 2015, and the field has grown rapidly since. These discoveries have been led by the LVK, an international collaboration of thousands of scientists. Sharing these discoveries has been a priority for the LVK. As a new field, communication has been essential to both publicise and explain results.
LVK activities span a range of audiences and media. Many resources are used in multiple contexts, such that utility can be amplified beyond the original target, e.g., paper plots used on social media, infographics in talks, and data audifications in musical performances.
Communicating LVK discoveries has a unique set of challenges, but each provides an opportunity:
  • GWs are an unfamiliar means of performing astronomy. Unlike traditional astronomy, we do not directly image our sources. However, GW data is suitable for audification, and a variety of tactile resources have been made. These make GW astronomy accessible to people with vision impairments, who are often excluded from accessible astronomy communication [42].
  • LVK results are produced by large teams. This makes it difficult for those outside to understand how discoveries are made, or to construct human-interest stories. However, this enables us to emphasise the importance of teamwork and international cooperation to science. The LVK’s breakthroughs are a direct contrast to the popular perception that scientific breakthroughs come from lone geniuses [43, 44]. The diversity of the LVK’s membership allows us to draw upon many different lived experiences, and demonstrate that scientists may come from any background.
  • After the initial first observations, many detections were similar to those discovered previously. The increase in the LVK’s detection rate is an experimental triumph—what was once extraordinary is now everyday. However, many detections are now not novel. It is therefore necessary to concentrate on the science enabled by a large set of detections. This has a benefit of reflecting that most scientific progress comes from careful study rather than a dramatic breakthrough.
While the LVK has benefitted from an existing interest in black holes and space, it has also faced hurdles, such as a tendency to misattribute discoveries to LIGO alone. As the field continues to mature, we expect that communication strategies will need to continue to evolve to reflect both the state of the field, and the interest and awareness of our audiences.
Acknowledgements
This material is based upon work supported by LIGO Laboratory which is a major facility fully funded by the US National Science Foundation and by the European Gravitational Observatory (EGO) which is funded by Centre National de la Recherche Scientifique (CNRS) in France, the Istituto Nazionale di Fisica Nucleare (INFN) in Italy, and the National Institute for Subatomic Physics (Nikhef) in the Netherlands. This paper has been assigned LIGO DCC number P2400039 (dcc.ligo.org/P2400039).
The Tactile Universe project was jointly funded by the University of Portsmouth and an STFC Public Engagement Legacy Award [ST/V001515/1]. CPLB acknowledges support of STFC grant ST/V005634/1. IC-C acknowledges support by the Spanish Agencia Estatal de Investigación through the Grant No. PID2021-125485NB-C21 funded by the MCIN/AEI/10.13039/501100011033, by the Generalitat Valenciana (Prometeo program for excellent research groups) through the Grant No. CIPROM/2022/49 and the Astrophysics and High Energy Physics program Grant No. ASFAE/2022/003 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1). HM is supported by the UK Space Agency Grant Nos. ST/Y004922/1; ST/V002813/1; ST/X002071/1.